OpenStudy (anonymous):

The volume of a cube is 125 cubic inches. What is the surface area of the cube?

OpenStudy (anonymous):

V = L * W * H in a cube the L, W, H are all the same so what number times itself 3 times is 125? The answer is 5 because 5 * 5 * 5 = 125. So L = 5, W = 5, H = 5 Surface area is the area it takes to cover the object. so You have 6 faces that are 5 x 5 so 6 * 25 = 150 in^2

OpenStudy (anonymous):

Thank you! I get this now!